OpenStudy (anonymous):

how do you solve 6x-10 over 7 divided by 9x-15 over 21?

OpenStudy (anonymous):

Well without it being equal to something you can't really 'solve' it...

OpenStudy (anonymous):

no there's a way to solve it because someone got the answer but i dont know what it is

OpenStudy (amistre64):

it can be reduced, simplified; but not really solved; its always gonna have an x in it unless that x factors out

OpenStudy (anonymous):

so what would the answer be and how would you get that answer?

OpenStudy (amistre64):

the other option is to plot the points on a graph that result from inputing values of x and the graph becomes the solution.... but without knowing tha tthis is a function with respect to x its rather redundant

OpenStudy (anonymous):

Is this the expression? \[(6x-10) \div {{9x - 15} \over 21}\]

OpenStudy (anonymous):

no.. 6x - 10 is over 7, but everything else is right

OpenStudy (anonymous):

\[{{6x-10} \over 7}\div {{9x - 15} \over 21}\]\[=2({{3x-5} \over 7}) \div 3({{3x-5} \over 7})\]\[={{2({{3x-5} \over 7})} \over {3({{3x-5} \over 7})}}\]\[ = {2 \over 3}\]

OpenStudy (anonymous):

Though you do need to stipulate that x cannot be equal to 5/3.

OpenStudy (anonymous):

one question.. how did you get both denominators 7?

OpenStudy (anonymous):

oh wait. I'm wrong.

OpenStudy (anonymous):

\[{{6x-10} \over 7}\div {{9x - 15} \over 21}\]\[=2({{3x-5} \over 7}) \div {3 \over 3}({{3x-5} \over 7})\]\[={{2({{3x-5} \over 7})} \over {1({{3x-5} \over 7})}} = 2,\ \forall x \ne {5 \over 3} \]

OpenStudy (anonymous):

Forgot I factored a 3 from the top and bottom of the fraction on the right.

OpenStudy (anonymous):

thx .can you help me with another problem?

OpenStudy (anonymous):

maybe

OpenStudy (anonymous):

Heres the problem...8-y over y squared + 4 minus y - 2 over y squared + 4

OpenStudy (anonymous):

It would be very helpful if you would write these up in the equation editor because there are a number of ways to interpret what you wrote there.

OpenStudy (anonymous):

oh i didn't know you could do that. because I'm kind of new to this website ....... 8-y/y squared + 4 minus y-2/y squared + 4? does that make a little more sense because I don't know how to do exponents so I just wrote it out

OpenStudy (anonymous):

\[{{8-y} \over y^2+ 4} - {{y-2} \over y^2+ 4} \] That?

OpenStudy (anonymous):

yah. how did u do that but anyway how do u solve it??

OpenStudy (anonymous):

Well they are over a common denominator already so just subtract the numerators.

OpenStudy (anonymous):

what answer would you get?